OpenStudy (anonymous):

Write the prime factorization of 112 (from least factor to greatest factor).

mathslover (mathslover):

Start it from 2 : 112 = 2 * 56 56 = 2 * 28 28 = 2 * 14 14= 2 * 7 That is you get : 112 = 2 * 2* 2 * 2* 7